OpenStudy (anonymous):

Dave rented a limousine for his wife's birthday. The hourly rate is $70. They used the limousine for 4 hours, plus Dave gave the driver a 10% tip. How much did he spend in total for the hourly charges plus tip? a $336 b $357 c $308 d $252

OpenStudy (anonymous):

@logan420

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute this: rent cost= 70*4=...? percentage = 0.1*(70*4)=... total cost= rent cost + percentage=...?

OpenStudy (anonymous):

@studyisland

OpenStudy (anonymous):

@bcniall_

OpenStudy (anonymous):

@aryandecoolest